2. A person with mass m= 90 kg is standing on bathroom scales in an elevator that has an acceleration a = 2 m/s2 upward. What is the reading on the scales? (Ans.: 1060 N).
Answer #1

Since acceleration is upwards

net force upward =mg+ma

so normal force =m(g+a)=1062 N

reading on scale =1062 N